Q1)On the basis of following information(All figures are in Rs),


A. CURRNT ACCOUNT


Exports= 150


Imports= 200


Services= 25


Income= ‐20


Transfer= 35


B. CAPITAL ACCOUNT


External Assistance(net)= 100


External Commercial Borrowing(net)= 150


Short Term Credit= ‐5


Banking Capital(net)= 250


Foreign Investment(net)= 300


Other Flows(net)= 70


Find,


 Trade Balance


 Current Account Balance


 Capital Account Balance

1
Expert's answer
2021-12-16T11:51:36-0500

Trade Balance=exports -imports

"=150-200=-50"

"Current\space account \space balance=Exports−Imports+Services+income+Transfers\\=Rs150−Rs200+Rs25−Rs20+Rs35\\=−Rs10"

"Capital \space account \space balance=Net \space external \space assistance+ Net \space external\space borrowing+Short−term \space credit+ Net\space banking \space capital+ Net \space foreign \space investment+ Other \space net \space flows=Rs100+Rs150−Rs5+Rs250+Rs300+Rs70\\=Rs865"